# Towards Flexible, Scalable, and Adaptive Multi-Modal Conditioned Face Synthesis

### 26 Dec 2023 v1

@@ Abstract
Recent progress in multi-modal conditioned face synthesis has enabled the creation of visually striking and accurately aligned facial images. Yet, current methods still face issues with scalability, limited flexibility, and a one-size-fits-all approach to control strength, not accounting for the differing levels of conditional entropy, a measure of unpredictability in data given some condition, across modalities. To address these challenges, we introduce a novel uni-modal training approach with modal surrogates, coupled with an entropy-aware modal-adaptive modulation, to support flexible, scalable, and scalable multi-modal conditioned face synthesis network. Our uni-modal training with modal surrogate that only leverage uni-modal data, use modal surrogate to decorate condition with modal-specific characteristic and serve as linker for inter-modal collaboration , fully learns each modality control in face synthesis process as well as inter-modal collaboration. The entropy-aware modal-adaptive modulation finely adjust diffusion noise according to modal-specific characteristics and given conditions, enabling well-informed step along denoising trajectory and ultimately leading to synthesis results of high fidelity and quality. Our framework improves multi-modal face synthesis under various conditions, surpassing current methods in image quality and fidelity, as demonstrated by our thorough experimental results.
Figure 1: Our method’s versatile synthesis capabilities, demonstrating high-fidelity facial image generation from a flexible combination of modalities including mask, text, sketch, lighting, expression, pose, and low-resolution images. Remarkably, these diverse face synthesis tasks are achieved within a single sampling process of a unified diffusion U-Net, demonstrating the method’s efficiency and the seamless integration of multi-modal information.
@@ 1 Introduction
Diffusion models have made remarkable achievements on a wide range of synthesis tasks in various domains e.g. image  [27, 16], audio [24, 17] ,video [1, 6] and motion [38, 35]. Within the sphere of image synthesis, the focus has increasingly shifted towards more controllable synthesis under multi-modal conditions [37, 21, 22, 20, 10]. These models can generate images that are not only visually compelling but also align well with given multimodal conditions. The emerging developments in this field are setting the stage for more dynamic and user-centric image synthesis, broadening its practicality and impact in real-world applications.
@@
One approach to achieving multi-modal conditioned synthesis, as explored in recent studies [22, 20], involves first developing uni-modal condition synthesis diffusion models and then averaging the noises generated by each modal branch, as shown in Figure 2 (a). This technique allows for flexible multi-modal conditions synthesis, enabling image generation under any combination of modal inputs. However, a notable limitation is its limited scalability. Each modality requires a distinct synthesis network, and the inference complexity increases linearly with the addition of each new modality. Building upon these foundations, Collaborative Diffusion [13] extends the work of [20, 22] by introducing a meta network to fuse the noise from each uni-modal branch, but require independently annotated multi-modal data for training. Another alternative methodology in the realm of multi-modal conditioned synthesis involves incorporating additional control mechanisms into basic synthesis models, as shown in Figure 2 (b). Notable examples include ControlNet [37] and T2I-adapter [21], which introduce layout control into text-conditioned image synthesis frameworks. These innovations enable image synthesis under the combined modal conditions of layout and text. However, this line of methods encounters limitations in terms of flexibility. They require specific tuning for each unique combination of modalities and require multi-modal annotated data. In addition, these methods often neglect the inherent differences in conditional entropy, a measure of unpredictability in data given some condition, associated with conditions from various modalities. As a result, they tend to assign uniform control strength across all modalities, disregarding their unique informational characteristics. In multi-modal synthesis scenarios, applying a fixed control strength uniformly across all modalities is suboptimal. This is particularly evident where conditions of high conditional entropy, under low control strength, fail to exert adequate influence on the image synthesis process. Conversely, an increased control strength might lead to over-fitting in cases involving conditions from modalities with low conditional entropy. Existing multi-modal conditioned synthesis [22, 20, 13] often assigns equal control across modalities, neglecting their distinct conditional entropy. As shown in Figure  5 where higher $w$ indicate higher control strength, a high-entropy modality (text), benefits from stronger control, while masks with lower entropy risk over-fitting.
@@
Figure 2: Core idea comparison between existing multi-modal synthesis approaches and our method. (a) Averaging noises from uni-modal diffusion models for flexible multi-modal synthesis, with scalability challenges due to linearly increased complexity. (b) Incorporation of additional control mechanisms in basic synthesis models for multi-modal synthesis conditioned synthesis, facing flexibility limitations. Each approach contends with the issue of uniform control strength across modalities with varying conditional entropy. (c) Our methods achieve multi-modal conditioned face synthesis within a single synthesis network, under flexible combination of conditions and dynamically adjust noise of diffusion step.
@@
To address the limitations of existing face synthesis frameworks, we introduce a Uni-modal Training approach with Modal Surrogate for enhanced flexibility and scalability, coupled with an entropy-aware modal-adaptive modulation mechanism for responsive adaptation to varying modal conditions. As shown in Figure 2, we set a distinct modal surrogate for each modality, designed to function as both a condition decorator for its respective modality and an inter-modal linker to facilitate collaboration between different modalities. During training, we only use uni-modal annotated data. The surrogate of the active modality is merged with the given condition, along with the surrogates of other modalities, to guide the de-noise process of a single diffusion U-Net. During this process, both the surrogate of the active modality and those of the other modalities are updated. The active modality’s surrogate acquires modality-specific information, while the others learn to interact with the active modality. This cycle is repeated for all considered modalities during training, ensuring each surrogate captures modality-specific context as a complement to the given conditions and serves as an interface for multi-modal conditioned face synthesis. Our synthesis network is highly flexible and scalable, capable of generating facial images under a variety of modal combinations within a single sampling process of a single diffusion model.
Given the disparate conditional entropy inherent to different modalities, we further develop an entropy-aware modal-adaptive modulation. This mechanism thoughtfully adjusts the noise levels, allowing the network to adapt its de-noising strategy to the distinct characteristics of each modality. Consequently, this adaptive approach ensures that each modality effectively influences the image generation process and the synthesis of facial images is faithful to all the given conditions.
@@
Using our novel approach of uni-modal training with modal surrogates, complemented by an entropy-aware, modal-adaptive modulation mechanism, our framework excels in generating high-fidelity facial images across multi-modal conditions. We extend beyond the standard modalities of masks, text and sketch ever considered in existing methods [13, 22, 20, 34, 37, 21], incorporating 3D face models and low-resolution images to enrich the multi-modal synthesis applications. This allows for a wide range of applications, e.g. uni-modal condition face synthesis including facial image super-resolution, lighting and sketch conditioned face synthesis, mask, sketch, text and lighting conditioned guided face generation, and so on, as shown in Figure  1.
@@
  • We propose an entropy-aware modal-adaptive modulation that dynamically adjusts the noise levels based on the conditional entropy of each modality. This allows our system to customize the denoising process, ensuring effective utilization of specific information from each modality and resulting in high-fidelity image synthesis that faithfully adheres to the given multi-modal conditions.
  • Our framework incorporates conditions of more modalities to significantly broaden the scope of our multi-modal synthesis capabilities, supporting a diverse range of face synthesis under uni-modal and complex multi-modal condition. This demonstrates the versatility and creative potential of our approach.
@@ 2.2 Conditioned Face Synthesis
In the evolving landscape of conditioned face synthesis, researchers have explored a range of controls to guide the synthesis process. This includes face generation under condition of sketches [33, 8], semantic masks [23], 3D face models [3, 2, 28, 5], low-resolution images [31], and textual descriptions [26, 30, 36, 19]. Recent methodologies [34, 22, 20, 13, 21, 37, 29, 12] have introduced multi-modal conditions into the face synthesis framework to achieve rich and flexible control than uni-modal condition. One line of methods [13, 22, 20] achieve multi-modal conditioned synthesis by firstly training uni-modal synthesis network in parallel and then fusing noise of each branch at each step.This technique allows for flexible synthesis across modalities. However, it tends to overlook the shared synthesis properties across different modalities, and suffer from an increased sampling complexity proportionate to the number of modalities. Another approach augments existing diffusion models with additional control mechanisms to handle multi-modal inputs [37, 21]. While this approach processes multiple modal inputs within a single network, it lacks flexibility and often requires additional training for new combinations of modalities. A third approach seeks to unify modal inputs by mapping them into a shared feature space, interpolating between various modality representations [29, 34]. However, this can sometimes overlook the inherent differences unique to each modality. Our method integrate the distinct characteristics of each modality and interaction among them within a unified, adaptive synthesis framework to support multi-modal conditioned face synthesis.
@@ 3 Method
Figure 3: Results of multi-modal training. The top row are conditions. The synthesis results are presented in the lower two rows. The resulting network can only synthesize pleasing results given all conditions and much of the guidance comes from the low-resolution image. The synthesis network tend to rely on modality of low condition entropy (LR) for synthesis and thus neglect modality with higher condition entropy.
Algorithm 1 Uni-modal Training with Modal Surrogate
@@
10:     $e^{n}\leftarrow e^{n}-\eta\frac{\partial L}{\partial\epsilon_{\theta}}\frac{\partial\epsilon_{\theta}}{\partial C^{m}}\frac{\partial C^{m}}{\partial e^{n}},n\neq m$
11:     $\triangleright$ Updating surrogate of other modality
12:     $e^{m}\leftarrow e^{m}-\eta\frac{\partial L}{\partial\epsilon_{\theta}}\frac{\partial\epsilon_{\theta}}{\partial C^{m}}\frac{\partial C^{m}}{\partial\tau^{m}}\frac{\partial\tau^{m}}{\partial e^{m}}$
13:     $\triangleright$ Updating surrogate of modality $m$
14:until converged
Figure 4: Uni-modal training with modal surrogate and entropy-aware modal-adaptive modulation mechanism. During training, we randomly sample uni-modal data, of which the condition is fused with its modal surrogate to learn modal-specific intrinsic and other modal surrogate to learn inter-modal collaboration. The fused features are sent to the diffusion U-Net to guide the de-noising process of the corrupted input image. The output noise is further modulated according to the condition features and UNet feature to adaptively adjust noise level given the conditions. The training process is provided in Algorithm  1.
Our goal is to synthesize facial image $x$ given conditions from multiple modalities, e.g. text and mask, which can be formally described as:
@@
where $M$ is the number of condition modalities. Recent works [22, 20] sample the multi-condition distribution in Equation 1 using stochastic gradient Langvein sampling and extend uni-modal classifier-free guidance for diffusion model [11] into multi-conditioned version:
As inllustrated in Figure  2 (a), each modality employs a separate synthesis network, ignoring shared aspects of image generation and causing sampling complexity to rise with more modalities. Control strengths $w_{m}$ remain the same across modalities, disregarding their unique entropy. Collaborative Diffusion[13] improves upon this by learning a noise-weighted network but still depends on individual networks and multi-modal data for training.
A concurrent technological trajectory [37, 21] is to achieve multi-modal conditioned synthesis by adding layout control to a pre-trained text-guide image synthesis framework, and can be described as:
where the $c_{l}$ is layout condition of a single modality and $w$ is actually control strength related only to textual modality. As shown in Figure  2 (b), such methodologies use a shared synthesis network parameter for multiple modalities. To add layout control of another modality, say $c_{3}$, this line of methods requires tedious respective fine-tuning for synthesis conditioned on other combinations like $c_{2}\&c_{3}$.
To alleviate the aforementioned issues, we instead formulate an efficient and flexible multimodal condition synthesis as shown in Figure  2 (c), which can be described as:
$\begin{split}\epsilon(x_{t},t,C)=(w+1)\epsilon_{\theta_{u},\theta_{w}}(x_{t},t,C)-\\ w(x_{t},t,C)\epsilon_{{\theta_{u},\theta_{w}}}(x_{t},t),\end{split}$ (4)
where $C$ represent any combination of modalities, e.g. $c_{1}$, $c_{1}\&c_{2}$ or $c_{1}\&c_{2}\&c_{3}$. $\theta_{u}$ and $\theta_{w}$ represent parameters of diffusion U-Net and weighting network in entropy-aware modal-adaptive modulation, respectively. We dynamically adjust the noise of diffusion net for different condition inputs. Note that we utilize a single model $\theta$ to handle conditions of different modalities with consistent sampling complexity. We describe in detail how we obtain the flexible and efficient synthesis framework with uni-modal training via modal surrogate in Section 3.1. In Section 3.2, we further elaborate our entropy-aware, modal-adaptive modulation mechanism to achieve multi-modal condition synthesis of high fidelity to given multi-modal conditions.
3.1 Uni-modal Training with Modal Surrogate
A straightforward approach for generating facial images under multi-modal conditions involves using multi-modally annotated data to train a synthesis network that aligns with multiple modalities. Such method may prioritize low-entropy modalities like low-resolution images, neglecting other modalities of higher entropy, as shown in figure 3. In multi-modal training, such shortcut issue arises when the network favors modalities with lower condition entropy to make learning process easier. However, this bias towards a specific modality might make the model less versatile across different modalities, not mentioning its requirements for multi-modal data. To fully learn how conditions guide the face synthesis process for each modality, we only utilize the condition of a single modality to train the synthesis network. Following [20, 22], the target in Equation  1 can be decomposed using Bayes’ theorem as
$\begin{split}P(x|c_{1},\dots,c_{M})=p(x)\prod_{m}{\frac{p(x|c_{m})}{p(x)}},\end{split}$ (5)
where uncondition distribution $p(x)$, uni-modal condition distribution $P(x|c_{m})$, and uni-modal conditional distribution $p(x|c_{1},\dots,c_{M})$ are all parameterized by a single network $\theta$ in our framework. To model multi-conditioned distribution $P(x|x_{1},\dots,x_{M})$ , we need to learn unconditional distribution $P(x)$, uni-modal conditional distribution $P(x|c_{m})$ of each modality. A key challenge is to learn inter-modal collaboration to achieve flexible multi-modal condition synthesis. To this end, we set for each modal $m$ a modal surrogate $e_{m}$ which serves as the modal decorator to represent modal-specific features as well as an interface to collaborate with different modalities. The presence of modal surrogate helps to capture the dependencies among modalities, which is critical when these modalities are actually not entirely independent. We rewrite Equation  5 as:
$\begin{split}P(x|c_{1},\dots,c_{M})=\\ P(x|e_{1},\dotsm e_{M})\prod_{m}{\frac{P(x|c_{m}+e_{m},e_{n}(n\neq m))}{P(x|e_{1},\dotsm e_{M})}}.\end{split}$ (6)
Alongside $c_{m}$ and $e_{m}$, $e_{n},n\neq m$ the surrogate of other modalities are also involved in training.
As shown in Figure  4, we first randomly choose a modality $m$ and here we take the modality mask as an example. The condition mask is sent to a mask encoder to extract condition feature $c_{m}$. The modal surrogate $e_{m}$ and condition feature $c_{m}$ are broadcast added together. In this way, $e_{m}$ decorates its control signal with modal-specific context. Other modality surrogates $e_{n},n\neq m$ are also fused with condition features for modality $m$, resulting in fused condition $C_{m}$. Here we instantiate the fuse process by broadcasting $e_{n},n\neq m$ and concatenating them with features for modality $m$ followed by several linear operations. The fused condition $C_{m}$ are sent to the diffusion U-Net to guide the denoising process of corrupted version $x_{t}$ of image $x_{0}$. The parameter of the diffusion U-Net and both the current training modality and other modal surrogates are updated. $e^{m}$ learns to decorate its signal with its modal-specific intrinsic and $e^{n},n\neq m$ learns to interact with the surrogate $e^{m}$ of the activate modality.
During each uni-modal training iteration for a specific modality $m$, the network not only focuses on the characteristics of modality $m$ but also learns how it relates to other modalities through their surrogates $e_{n},n\neq m$. The presence of other modalities’ surrogate enables the network to grasp the contextual relationships that might exist between different modalities. This strategy ensures that the network is not only trained on the specifics of each modality but also understands how different modalities interact, ultimately leading to a model capable of synthesizing outputs from complex multi-modal inputs. By learning cross-modal interactions in this manner, there is no need to design separate training processes for every possible combination of modalities. This approach prevents the network from becoming overly reliant on some modality that might have lower condition entropy, thereby maintaining versatility across different types of data inputs.
Condition
$w=0$
$w=2$
$w=9$
She is in the thirties. No beard.
Condition
$w=0$
$w=2$
$w=9$
$w=20$
Figure 5: Uni-modal synthesis results given different control strength ($w$). To generate facial images of high fidelity and quality, text guided face synthesis require higher control strength than mask.

### 21 Mar 2024 v2

# https://jingjingrenabc.github.io/multimodal-face-synthesis/

@@ Abstract
Recent progress in multi-modal conditioned face synthesis has enabled the creation of visually striking and accurately aligned facial images. Yet, current methods still face issues with poor scalability, limited flexibility, and a one-size-fits-all approach to control strength for various conditions. To address these challenges, we introduce a novel uni-modal training approach with modal surrogates, coupled with an entropy-aware modal-adaptive modulation, to support flexible, scalable, and adaptive multi-modal conditioned face synthesis. Our modal surrogate decorate condition with modal-specific characteristic and serve as linker for inter-modal collaboration, resulting in a highly scalable and flexible multi-modal face synthesis framework. The entropy-aware modal-adaptive modulation finely adjust diffusion noise according to modal-specific characteristics and given conditions. It enables well-informed step along de-noising trajectory and ultimately leads to synthesis results of high fidelity. Building upon our scalable, flexible and adaptive multi-modal synthesis framework, we efficiently incorporate more modalities and support a wide range of face synthesis applications. Our extensive experiments demonstrate our method’s superiority for multi-modal face synthesis.
Keywords:
Multi-modal face synthesis Diffusion model
Figure 1: Our method’s versatile synthesis capabilities, demonstrating high-fidelity facial image generation from a flexible combination of modalities. Remarkably, these diverse face synthesis tasks are achieved within a single sampling process of a unified diffusion U-Net, demonstrating the method’s efficiency and the seamless integration of multi-modal information.
@@ 1 Introduction
Diffusion models have made remarkable achievements on a wide range of synthesis tasks in various domains e.g. image  [29, 17], audio [25, 18] ,video [1, 6] and motion [41, 38]. Within the sphere of image synthesis, the focus has increasingly shifted towards more controllable synthesis under multi-modal conditions [40, 22, 23, 21, 10, 42, 26]. These models can generate images that are not only visually compelling but also align well with given multi-modal conditions. The emerging developments in this field are setting the stage for more dynamic and user-centric image synthesis, broadening its practicality and impact in real-world applications.
@@
One approach to achieve multi-modal conditioned synthesis, as explored in recent studies [23, 21, 13], involves first developing uni-modal condition synthesis diffusion models and then fusing the noises generated by each modal branch, as shown in Fig. 2 (a). This technique allows for flexible multi-modal conditions synthesis, enabling image generation under any combination of modal inputs. However, a notable limitation is its limited scalability. Each modality requires a distinct synthesis network, and the inference complexity increases linearly with the involvement of new modality. Another alternative approach [26, 22, 40, 42] to multi-modal conditioned synthesis involves incorporating additional control mechanisms into basic synthesis models, as shown in Fig. 2 (b). These innovations enable image synthesis under the combined modal conditions of layouts and text. However, these methods fall short of flexibility. They rely on specific tuning for each unique combination of modalities and require multi-modal annotated data.
Furthermore, conditions from different modality naturally exhibits diverse condition entropy, a measure of unpredictability in data given some condition. A condition of higher entropy requires higher control strength to exert adequate influence on face synthesis, while some condition of lower entropy suffer from over-fitting with the same control strength. As shown in Fig. 3 where higher $w$ indicate higher control strength, a high-entropy modality (text), benefits from stronger control, while masks with lower entropy risk over-fitting. Existing multi-modal conditioned synthesis [23, 21, 13, 42, 26] often neglect such inherent differences in conditional entropy for various modalities. They consequently assign equal control across modalities, and produce results of poor alignment with given multi-modal condition.
@@
Figure 2: Core idea comparison between existing multi-modal synthesis approaches and our method. (a) Fusing noises from multiple uni-modal diffusion models. (b) Incorporation of additional control mechanisms in basic synthesis models for multi-modal synthesis conditioned synthesis. (c) Our method achieve multi-modal conditioned face synthesis within a single synthesis network, under flexible combination of conditions and dynamically adjust noise of diffusion step.
@@
To address the limitations of existing methods in terms of scalability, flexibility and adaptivity, we introduce a uni-modal training approach with modal surrogate for enhanced flexibility and scalability, coupled with an entropy-aware modal-adaptive modulation mechanism for responsive adaptation to varying modal conditions. As shown in Fig. 2, we set a distinct modal surrogate for each modality, designed to function as both a condition decorator for its respective modality and an inter-modal linker to facilitate collaboration between different modalities. During training, we only use uni-modal annotated data. The surrogate of the active modality is merged with the given condition, capturing modality-specific context as a complement to the given conditions. On the other hand, surrogates of other modalities is also involved with the given condition, to learn collaboration with the activate modality. With the surrogate decorating function and inter-modal learning, the resulting network can discriminate and process condition among different modalities. Therefore our synthesis network is highly flexible and scalable, capable of generating facial images under a variety of modal combinations within a single sampling process of a single diffusion model. To take into full consideration the disparate conditional entropy inherent to different modalities, we further develop an entropy-aware modal-adaptive modulation mechanism. By thoughtfully adjusting the noise levels, this mechanism allows the network to adapt its de-noising strategy to the distinct characteristics of each modality. Consequently, this adaptive approach ensures that each modality sufficiently guides the image generation process, leading to faithful synthesis to all given conditions.
Figure 3: Uni-modal synthesis results given different control strength $w$. To generate facial images of high fidelity and quality, text and mask require different control strength due to their entropy difference.
@@
Using our novel approach of uni-modal training with modal surrogates, equipped with by an entropy-aware, modal-adaptive modulation mechanism, our framework excels in generating high-fidelity facial images across flexible combination of conditions. We extend beyond the standard modalities of masks, text and sketch ever considered in existing methods [13, 23, 21, 37, 40, 22] for face synthesis, incorporating lighting, pose, expression and low-resolution images to enrich the multi-modal synthesis applications. This allows for a wide range of applications, e.g., various uni-modal face synthesis in Fig. 1 (a), pose and text conditioned face synthesis in Fig. 1 (b), lighting and sketch conditioned face synthesis in Fig. 1 (d), and so on.
@@
  • We propose an entropy-aware modal-adaptive modulation that dynamically adjusts the noise levels based on the conditional entropy of each modality. This allows our system to finely adjust the de-noising process, ensuring effective utilization of specific information from each modality and resulting in face synthesis of high fidelity and quality.
  • Our framework incorporates conditions of more modalities to significantly broaden the scope of our multi-modal face synthesis capabilities, supporting a diverse range of face synthesis under uni-modal and complex multi-modal condition. This demonstrates the versatility and creative potential of our approach.
@@ 2.2 Conditioned Face Synthesis
In the evolving landscape of conditioned face synthesis, researchers have explored a range of controls to guide the synthesis process. This includes face generation under condition of sketches [36, 8], semantic masks [24], 3D face models [2, 3, 30, 5], low-resolution images [35], and textual descriptions [28, 32, 39, 20]. Recent methodologies [37, 23, 21, 13, 22, 40, 31, 12] have introduced multi-modal conditions into the face synthesis framework to achieve rich and flexible control than uni-modal condition.
One line of methods [13, 23, 21] sample multi-condition distributions (Eq. 3) using stochastic gradient Langvein sampling, converting uni-modal classifier-free guidance (CFG)[11] for diffusion model into a multi-conditioned version:
where $C=\{c_{1},...,c_{M}\}$ respresents multi-modal condition and $w_{m}$ denotes the controlling strength of the $m$-th modality. As defined in Eq. 1, each modality $m$ employs an individual synthesis network, ignoring the shared aspects of image generation. This technique allows for flexible synthesis across modalities. However, it tends to overlook the shared synthesis properties across different modalities, and suffer from an increased sampling complexity proportionate to the number of modalities.
A concurrent technological trajectory [40, 22, 42, 26] achieve multi-modal conditioned synthesis by adding layout control to a pre-trained text-guide image synthesis model, which can be presented as:
where $C_{l}$ is the layout condition of a single or multiple modalities. $w$ represents the controlling strength of the textual modality. As shown in Fig. 2 (b), these methods use a parameter-sharing network for multiple modalities guided synthesis. However, they directly integrate controlling features of different modalities for multiple layout conditioned synthesis, tending to produce inferior results due to the lack of inter-modal collaboration during the model learning. While this approach processes multiple modal inputs within a single network, it lacks flexibility and often requires additional training for each combinations of modalities. Our method integrate the distinct characteristics of each modality and interaction among them within a unified, adaptive synthesis framework to support multi-modal conditioned face synthesis.
@@ 3 Method
Figure 4: Results of multi-modal training. The left are input multi-modal conditions. The synthesis results are presented in the right part. The resulting network can only synthesize pleasing results given all conditions and much of the guidance comes from the low-resolution image. The synthesis network tend to rely on modality of low condition entropy (LR) for synthesis and thus neglect modality with higher condition entropy.
Algorithm 1 Uni-modal Training with Modal Surrogates
@@
10:    $e^{n}\leftarrow e^{n}-\eta\frac{\partial L}{\partial\epsilon_{\theta}}\frac{\partial\epsilon_{\theta}}{\partial C^{m}}\frac{\partial C^{m}}{\partial e^{n}}$
11:    $\triangleright$ Updating surrogate of ALL modality
12:until converged
Figure 5: Uni-modal training with modal surrogates and entropy-aware modal-adaptive modulation mechanism. During training, we randomly sample uni-modal data, of which the condition is fused with its modal surrogate to learn modal-specific intrinsic and other modal surrogate to learn inter-modal collaboration. The fused features are sent to the diffusion U-Net to guide the de-noising process of the corrupted input image. The output noise is further modulated according to the condition features and UNet feature to adaptively adjust noise level given the conditions. The training process is provided in Algorithm 1.
The goal to synthesize facial image $x$ given conditions from multiple modalities, e.g., text and mask, can be formally formulated as learning:
@@
where $c_{m}$ denotes the $m$-th condition modalities and $M$ is the number of condition modalities. we devise an efficient, flexible, and adaptive multi-modal conditioned face synthesis, as shown in Fig. 2(c). In the following sections, we first elaborate in detail our uni-modal training via modal surrogates for efficient and flexible synthesis (Sec. 3.1). Afterwards, we present our adaptive weighting mechanism for adaptive synthesis (Sec. 3.2).
3.1 Uni-modal Training with Modal Surrogates
To ensure efficient synthesis, we adopt a shared synthesis network for conditions from all modalities as previous works [22, 40, 26, 42]. This network is required to discriminate between those different modalities, to support various conditioned face synthesis within a single network. To this end, we assign a modal surrogate to each modality, for decorating the condition with its modal-specific information. The learning objective can be represented as
$L:=\mathbb{E}_{t,x^{i}_{0},\epsilon\sim\mathcal{N}(0,1)}[||\epsilon_{\theta}(x^{i}_{t},f(c^{i}_{m},e_{M}),t),\epsilon||_{2}],$ (4)
where $x^{i}_{t}$ is noised version of the $i$-th input image $x^{i}_{0}$. $f(\cdot)$ denotes the fusion operation of condition $c^{i}_{m}$ and the modal surrogate $e_{m}$ of its modality. The modal surrogate is a learnable token. We instantiate $f(\cdot)$ by broadcasting the modal surrogate and adding it to the condition $c_{m}$. The modal surrogate is to decorate condition with modal intrinsic and helps network discriminate among different modal. Thus far, our model is capable of synthesizing faces given various single modal condition, e.g., mask, sketch or text.
To achieve multi-modal face synthesis, a straightforward way is to train a unified model with paired multi-modal annotated data (i.e., $(x^{i},c_{1}^{i},...c_{M}^{i})$) to generate faces that align with the multi-modal conditions simultaneously. However, it would be extremely complex and tedious to consider all those combinations during training as the total number of combinations of $M$ modalities could be $C_{M}^{1}+C_{M}^{2}+...+C_{M}^{M}$. Moreover, in multi-modal training, the network may prioritize the most informative condition to loosen the learning difficulty. To demonstrate this, we illustrate a representative example where network relies on the low resolution image condition most to render the face, while neglecting other modalities in Fig. 4. To solve the above issue, we design a inter-modal learning mechanism that enables the network to not only learn form uni-modal data $x^{i},x^{i}_{m}$ as in Eq. 4, but also take fully advantage of rich modality cues from modal surrogates of other modalities. Therefore, the training objective in Eq. 4 can be then reformulated as:
$L:=\mathbb{E}_{t,x_{0},\epsilon\sim\mathcal{N}(0,1)}[||\epsilon_{\theta}(x^{i}_{t},f(c^{i}_{m},e_{1},...,e_{M}),t),\epsilon||_{2}],$ (5)
where the modal surrogates of other modalities are also involved when data of condition $c_{m}$ is fed for training. Note that the modal surrogates of modal $c_{m}$ in Eq. 5 is similarly defined as in Eq. 4, but it additionally learns inter-modal interaction with other modals. For modal surrogates of other modals $n$ ($n\neq m$), their modal surrogates also get updated. In this way the surrogates of some modal also learn from data of other modalities, and thus effectively capture inter-modal interactions with modal $m$ during the whole training process. Such inter-modal modelling helps aid network in grasping inter-modal collaboration via uni-modal training only. Consequently our method can perform flexible synthesis given arbitrary combination of condition modalities within a single sampling process of a single synthesis network.